### v2

@@ 3.5 Theoretical Analysis
[118] evaluated the performance of self-supervised learning (SSL) by determining whether the trained model represents enough information to obtain the data distribution, given additional information about the family distribution. The SSL task chosen for this assessment was the masked prediction task.
@@
[120] presented a Bayesian theoretical view of the underlying mechanisms of masked image modeling. They hypothesized that masked pre-training, and implicitly the reconstruction error minimization, can be equivalent to maximizing the marginal likelihood, demonstrating this proposition on language and vision models.
@@ 3.6 Model Architecture
Figure 16: MCMAE architecture, courtesy of [128] (image licensed under CC BY 4.0). MCMAE falls into the category of reconstruction-based methods illustrated in Figure 2, as it leverages an ${L}^{2}$ loss between the masked patches and the corresponding patches from the input image.
Motivated by leveraging the masked pre-training strategy of autoencoders with convolutional layers, [128] presented an architecture that combines them, as illustrated in Figure 16. The encoder consists of three stages, where the first two are convolutional and the last one is transformer-based. First, a mask is sampled to determine which tokens are visible. The mask is then upsampled at the resolutions of the first two stages, in order to be used by masked convolutional blocks. Information from the first two stages is added to the resulting tokens of the third stage, and then, they are linearly projected. Finally, all tokens (predicted and masked) are decoded into the original pixel space. The authors state that the advantage of this method is represented by the multi-scale features learned by the encoder. Having the same goal of adopting convolutional layers in MIM pipelines, [126] presented a MIM method that leverage CNNs. As in the MAE framework, the input is split in non-overlapping patches and some of them are masked. The encoder is composed of sparse convolutional layers, thus preserving the mask pattern intact along the feature maps. The decoder, used for reconstructing the original image, is composed of upsampling blocks that receive the previous layer, as well as the encoder’s features on the same level. The masking regions are replaced with a mask embedding.
@@
3.7 Target Features and Objective Function
We next discuss studies that make contributions to both the target features and the objective function. While most approaches use a deep encoder, either frozen [132, 133] or updated online [21, 134, 135, 136, 137] during training, to extract the target features, there are also methods that extract their features directly from the raw input signal [138], eliminating the need for a neural network. In terms of objective functions, most of the works focus on integrating a contrastive loss component [21, 135, 136], but there are also a few methods that leverage a multi-scale reconstruction loss [138].
@@ 3.7 Target Features and Objective Function
3.8 Model Architecture and Objective Function
A number of papers contributed to both the model architecture and the objective function. An important subset of these studies made their architectural changes as a result of integrating the new objective [142, 76, 143, 144]. For example, the introduction of adversarial loss by [142] requires an additional discriminator and the multi-scale reconstruction loss used by [76] implies additional adapters to ensure compatibility between features. Another direction of research that requires both architectural and objective function modifications is the integration of multiple modalities [145, 6]. In such scenarios, each modality typically requires its own dedicated modules, leading to architectural adjustments, while an additional loss term is introduced to ensure alignment between the features of the two modalities.
@@ 3.8 Model Architecture and Objective Function
Inspired by masked autoencoders that process both visual and textual modalities, [146] adapted the masking pre-training framework for optical and radar inputs. Both modalities, after being aligned, tokenized and randomly masked, are encoded using an individual encoder, and then they are jointly processed by a multimodal encoder. The resulting embedding is decoded to reconstruct both input images. A contrastive loss between the mean embeddings of the individual encoders is adopted to match sensor measurements from the same timestamp, while maximizing the difference between those at different timestamps.
@@
3.9 Masking Strategy and Target Features
@@
Other works focused on implementing MIM for point clouds. On the one hand, [149] presented I2P-MAE, a method designed to learn better 3D features by reconstructing masked point clouds. The approach leverages 2D pre-trained models to keep the important point tokens visible while masking. Moreover, the 2D models are used to get the target representations for a semantic reconstruction loss that is applied on the visible tokens, after the decoder. On the other hand, some works harnessed 2D images in their point-cloud methods. [155] proposed a pre-training method for 3D point clouds by leveraging the corresponding 2D visual representation. The 3D points and their 2D projections are jointly encoded. The resulting encoding is randomly masked and passed through a two-stage decoder. First, there is a shared decoder, then each representation continues through a separate module to reconstruct both input modalities. [152] utilized the mask-reconstruction strategy in 3D segmentation due to the lack of supervised data, as well as the domain difference between training and testing data. During training, given a pair composed of a 2D image and a 3D point cloud, patches from one modality are masked and the model tries to estimate them using the other modality. A CNN backbone with a lower masking ratio is employed. Having two different datasets (source-labeled and target-unlabeled), the MIM is performed on both datasets, while the supervised task is performed only on the former.
@@
Inspired by MAE, [153] modified the MAE framework in order to boost performance when applied to downstream classification tasks in few-shot scenarios. Rather than reconstructing the original pixel space, the authors proposed to operate in the latent representation space of a frozen backbone. Two subsets of images, called support and query, are firstly embedded, the latter’s embeddings being masked. Then, the support embeddings are encoded, concatenated with mask tokens, and decoded to estimate the corresponding query embeddings, MSE being the computed reconstruction loss. After being pre-trained with a large labeled dataset, the method is further trained on the smaller dataset (containing few examples per class) in order to learn the global information about each class.
3.10 Masking Strategy and Model Architecture
@@
Harnessing multiple sources of information during MIM pre-training was demonstrated to bring great benefits, especially for multimodal downstream tasks. [26] proposed an MAE framework that can be used with multiple modalities. Given a set of modalities, each is tokenized into a common representation form. Rather than using all tokens, only two subsets from each modality are sampled: one that is encoded and the other one that is masked and then reconstructed. A joint encoder is employed for all input types. Still, a modality embedding is added to each token to indicate its source. In the cross-attention layers of the decoder, the embeddings corresponding to one modality are masked from the rest, while using all resulting tokens from the encoder. Besides demonstrating good results on downstream tasks, the method shows promising cross-modal generative capabilities. To leverage the information from multiple input data types and learn richer feature representations, [160] proposed to pre-train a network with multiple modalities. Their self-supervised method involved three types of data: an RGB image, a depth map and a semantic segmentation map. Using the ViT architecture, the modalities are split into patches and projected into tokens (with separate layers for each modality). Then, a great proportion of tokens from each modality is masked, while the rest (unmasked) are encoded (using a joint encoder) and concatenated with the masked tokens. A separate decoder for each data type is used to reconstruct the corresponding input. In the first stage of the decoder composed of a cross-attention layer, the tokens associated with the respective modality are given as queries, while the keys and values given are from all tokens. Experiments carried out on downstream tasks for all three data types indicate that the proposed strategy shows competitive results. PiMAE [159] is a self-supervised framework based on MAE, which learns representations that capture interactions between point clouds and images. Overall, the approach is based on the usual reconstruction objective for each modality. However, in contrast to MAE, the masking strategy in this case is designed to be complementary between the two modalities. In terms of architectural changes, the encoder and decoder include some common blocks between the two modalities, but they also have modality specific layers. [157] combined two sources of information (Hematoxylin and Eosin and Immunohistochemical staining images) to detect breast cancer, adopting MAE as the base for their method. Both images are split into patches, some of which are randomly masked, while the remaining visible patches are fed together through a ViT-based model. The resulting embeddings, together with the learnable mask embeddings, are further processed by two self-attention modules (one specific for each modality), as well as a cross-attention module (that processes all embeddings). Finally, two separate decoders reconstruct the original images, each receiving the modality-specific embeddings, as well as the inter-modal ones. More recently, [163] employed the MAE framework to learn robust cross-modal representations. Their primary contribution is the MultiModal 3D Interaction module, which takes as input the concatenation of masked 3D volume features from both video frames and LiDAR sensors. Built upon self-attention layers and feed-forward networks, this module outputs a representation that is then split along the channel dimension and fed into modality-specific decoders to reconstruct the original inputs.
@@
3.11 Masking Strategy and Objective Function
A body of works have contributed to both the masking strategy and the objective function. These works either introduced a novel masking strategy that excels in hiding the salient information [1, 165, 33], or adapted the masking strategy to a different input type [105, 166]. Further, to maximize the benefit of using a new masking strategy, the objective function is altered to be more suitable for recovering the original input source [105, 32, 33, 167]. For example, [105] performed the masking in the discrete space of a discrete Variational Autoencoder (dVAE) and, due to this change, the reconstruction loss is the cross-entropy. Another example is the work of [166], in which the contrastive loss is employed to align the token representations of two different masked versions of the same image.
@@ 3.11 Masking Strategy and Objective Function
[167] slightly modified the MAE model by adding Gaussian noise to all pixels in the input image. With this change, besides reconstructing the masked regions, the objective is further extended to decode all denoised patches. The experiments demonstrated the superiority of this method over the original MAE. [33] introduced some improvements to the teacher-student MIM based on contrastive pre-training. The first addition is a ranking component that divides the patches into two: the ones that contain salient information and the meaningless ones. The former subset is passed to the student, while the latter is given to the teacher, both being masked and leaving the rest visible. The first objective is to reconstruct the masked patches of the student, as they are harder to predict due to containing more salient information. A second loss is used to align the globally encoded representations of the two subsets by maximizing the similarity between the embeddings of the [CLS] token. While the same encoder is used for both models, the gradients do not flow through the teacher.
@@
[5] introduced a masking strategy that selects important patches according to their similarity with a given text, computed in the CLIP embedding space. They further employ a contrastive loss to align text features with the features extracted from the unmasked patches.
@@
A couple of papers opted for a contrastive objective between two different views of an image obtained by distinct augmentation. Taking a different masking SSL approach, [168] proposed a method whose objective is to be more robust for downstream tasks. On the input image, two different sets of augmentations are applied. The augmented images are encoded, and a projection layer transforms the encoded representations into multivariate Gaussian distributions. Finally, a series of learnable masks are applied. The difference between both masked probabilistic embeddings is computed, the objective being that of minimizing this difference. [3] presented a pre-training framework that combines MIM and contrastive learning. Different subsets (a more aggressive one and a lighter one) of augmentations are applied to the input image to create two views, the latter having a portion of its tokens masked. Then, two vision transformer models (where the one for the unmasked image is an EMA of the other model) are used to encode the tokenized patches. The contrastive loss is applied only between the corresponding corrupted patches.
3.12 Masking Strategy and Downstream Task
@@
A few papers adopted MIM for increasing the quality of images. The Saturated Mask AutoEncoder (SMAE), introduced by [175], is a two-stage approach for few-shot HDR imaging. The first stage focuses on representation learning, which is performed via masked image modeling. In this stage, the method creates two additional images from the original frame using exposure adjustment. Next, all three images are randomly masked with a high masking ratio before passing them to the model. LEMaRT [173] is an effective pre-training framework when applied on image harmonization as a downstream task. In this approach, the masked patches are replaced with the patches taken from a perturbed version of the original image. The research also investigated what is the best strategy for creating the masks, concluding that random masking works best for image harmonization. Magnetic Resonance scans are generated in k-space and then transformed in the image domain with the inverse Fourier transform. To obtain an image with high quality and fidelity, the k-space needs to be fully sampled, but this is not realistic in most scenarios. To this end, [176] proposed to leverage the MAE framework by masking data in the k-space, represented in 3D: height, width and time, along the first dimension. By reconstructing the missing tokens via the $L^{1}$ loss, the adopted ViT models learn a rich feature representation that is able to estimate the unsampled data from k-space at inference. The estimated k-space is further refined using three sequential transformer-based decoders (one for each pair of dimensions), and employing the High-Dynamic Range loss after each decoder.
@@
3.13 Target Features and Downstream Task
@@
3.14 Model Architecture and Target Features
@@
3.15 Objective function, Downstream Task and Theoretical Analysis
@@
3.16 Masking Strategy, Model Architecture and Objective Function
@@
3.17 Masking Strategy, Target Features and Objective Function
@@
3.18 Masking Strategy, Model Architecture, Downstream Task and Objective Function

### v3

@@ 3.5 Theoretical Analysis
[135] evaluated the performance of self-supervised learning (SSL) by determining whether the trained model represents enough information to obtain the data distribution, given additional information about the distribution family. The SSL task chosen for this assessment was the masked prediction task.
@@
Figure 15: MCMAE [140] falls into the category of reconstruction-based methods illustrated in Figure 2, as it leverages an ${L}^{2}$ loss between the masked patches and the corresponding patches from the input image.
@@ 3.6 Model Architecture
Motivated by leveraging the masked pre-training strategy of autoencoders with convolutional layers, [140] presented an architecture that combines them, as illustrated in Figure 15. The encoder consists of three stages, where the first two are convolutional and the last one is transformer-based. First, a mask is sampled to determine which tokens are visible. The mask is then upsampled at the resolutions of the first two stages, in order to be used by masked convolutional blocks. Information from the first two stages is added to the resulting tokens of the third stage, and then, they are linearly projected. Finally, all tokens (predicted and masked) are decoded into the original pixel space. The authors state that the advantage of this method is represented by the multi-scale features learned by the encoder. Having the same goal of adopting convolutional layers in MIM pipelines, [144] presented a MIM method that leverages CNNs. As in the MAE framework, the input is split into non-overlapping patches and some of them are masked. The encoder is composed of sparse convolutional layers, thus preserving the mask pattern intact along the feature maps. The decoder, used for reconstructing the original image, is composed of upsampling blocks that receive the previous layer, as well as the encoder’s features on the same level. The masking regions are replaced with a mask embedding.
@@
3.7 Features and Objective
We next discuss studies that make contributions to both the target features and the objective function. While most approaches use a deep encoder, either frozen [149, 150] or updated online [21, 151, 152, 153, 154] during training, to extract the target features, there are also methods that extract their features directly from the raw input signal [155], eliminating the need for a neural network. In terms of objective functions, most of the works focus on integrating a contrastive loss component [21, 152, 153], but there are also a few methods that leverage a multi-scale reconstruction loss [155]. Another important direction is closely following the strategy in BERT [18], using discrete tokens and predicting the id from a learned vocabulary [85, 86]. These tokens are obtained using a variational autoencoder (VAE), the pixels being mapped to a discrete latent space [156]. With a similar goal in mind, other works map each token into the nearest entry from a codebook of learned embeddings [157, 34].
@@ 3.7 Features and Objective
[85] adapted the pre-training strategy from BERT to images. First, the image patches are mapped to discrete visual tokens using a discrete Variational Autoencoder (dVAE) [156]. Then, some tokenized patches are randomly masked and replaced with a special mask token. These are then passed through the encoder together with the visible tokens. The final objective is to predict the id of the masked tokens from the learned vocabulary of visual tokens. Instead of a decoder, this method uses only a classification head, which is dropped during fine-tuning. [160] extended BEiT by employing multiple tokens for each patch. Inspired by the aforementioned work, [86] proposed to learn both the tokenizer and the encoder. Their method adopts the teacher-student framework, such that the student tries to reconstruct the corresponding masked tokenized patches from the teacher and to match the [CLS] token.
3.8 Architecture and Objective
A number of papers contributed to both the model architecture and the objective function. An important subset of these studies made their architectural changes as a result of integrating the new objective [161, 94, 162, 163]. For example, the introduction of adversarial loss by [161] requires an additional discriminator, and the multi-scale reconstruction loss used by [94] implies additional adapters to ensure compatibility between features. Another direction of research that requires both architectural and objective function modifications is the integration of multiple modalities [164, 6]. In such scenarios, each modality typically requires its own dedicated modules, leading to architectural adjustments, while an additional loss term is introduced to ensure alignment between the features of the two modalities.
@@ 3.8 Architecture and Objective
Inspired by masked autoencoders that process both visual and textual modalities, [165] adapted the masking pre-training framework for optical and radar inputs. Both modalities, after being aligned, tokenized and randomly masked, are encoded using an individual encoder, and then, they are jointly processed by a multimodal encoder. The resulting embedding is decoded to reconstruct both input images. A contrastive loss between the mean embeddings of the individual encoders is adopted to match sensor measurements from the same timestamp, while maximizing the difference between those at different timestamps.
@@
3.9 Masking Strategy and Features
@@
Other works focused on implementing MIM for point clouds. On the one hand, [168] presented I2P-MAE, a method designed to learn better 3D features by reconstructing masked point clouds. The approach leverages 2D pre-trained models to keep the important point tokens visible while masking. Moreover, the 2D models are used to get the target representations for a semantic reconstruction loss that is applied on the visible tokens, after the decoder. On the other hand, some works harnessed 2D images in their point-cloud methods. [174] proposed a pre-training method for 3D point clouds by leveraging the corresponding 2D visual representation. The 3D points and their 2D projections are jointly encoded. The resulting encoding is randomly masked and passed through a two-stage decoder. First, there is a shared decoder, then each representation continues through a separate module to reconstruct both input modalities. [171] utilized the mask reconstruction strategy in 3D segmentation due to the lack of supervised data, as well as the domain difference between training and testing data. During training, given a pair composed of a 2D image and a 3D point cloud, patches from one modality are masked and the model tries to estimate them using the other modality. A CNN backbone with a lower masking ratio is employed. Having two different datasets (source-labeled and target-unlabeled), the MIM is performed on both datasets, while the supervised task is performed only on the former.
@@
Inspired by MAE, [172] modified the MAE framework in order to boost performance when applied to downstream classification tasks in few-shot scenarios. Rather than reconstructing the original pixel space, the authors proposed to operate in the latent representation space of a frozen backbone. Two subsets of images, called support and query, are firstly embedded, the latter’s embeddings being masked. Then, the support embeddings are encoded, concatenated with mask tokens, and decoded to estimate the corresponding query embeddings, MSE being the computed reconstruction loss. After being pre-trained with a large labeled dataset, the method is further trained on a smaller dataset (containing few examples per class) in order to learn global information about each class.
3.10 Masking Strategy and Architecture
@@
Harnessing multiple sources of information during MIM pre-training was demonstrated to bring great benefits, especially for multimodal downstream tasks. [26] proposed an MAE framework that can be used with multiple modalities. Given a set of modalities, each is tokenized into a common representation form. Rather than using all tokens, only two subsets from each modality are sampled: one that is encoded and the other one that is masked and then reconstructed. A joint encoder is employed for all input types. Still, a modality embedding is added to each token to indicate its source. In the cross-attention layers of the decoder, the embeddings corresponding to one modality are masked from the rest, while using all resulting tokens from the encoder. Besides demonstrating good results on downstream tasks, the method shows promising cross-modal generative capabilities. To leverage the information from multiple input data types and learn richer feature representations, [179] proposed to pre-train a network with multiple modalities. Their self-supervised method involved three types of data: an RGB image, a depth map and a semantic segmentation map. Using the ViT architecture, the modalities are split into patches and projected into tokens (with separate layers for each modality). Then, a great proportion of tokens from each modality is masked, while the rest (unmasked) are encoded (using a joint encoder) and concatenated with the masked tokens. A separate decoder for each data type is used to reconstruct the corresponding input. In the first stage of the decoder composed of a cross-attention layer, the tokens associated with the respective modality are given as queries, while the given keys and values are from all tokens. Experiments carried out on downstream tasks for all three data types indicate that the proposed strategy shows competitive results. PiMAE [178] is a self-supervised framework based on MAE, which learns representations that capture interactions between point clouds and images. Overall, the approach is based on the usual reconstruction objective for each modality. However, in contrast to MAE, the masking strategy in this case is designed to be complementary between the two modalities. In terms of architectural changes, the encoder and decoder include some common blocks between the two modalities, but they also have modality specific layers. [176] combined two sources of information (Hematoxylin and Eosin and Immunohistochemical staining images) to detect breast cancer, adopting MAE as the base for their method. Both images are split into patches, some of which are randomly masked, while the remaining visible patches are fed together through a ViT-based model. The resulting embeddings, together with the learnable mask embeddings, are further processed by two self-attention modules (one specific for each modality), as well as a cross-attention module (that processes all embeddings). Finally, two separate decoders reconstruct the original images, each receiving the modality-specific embeddings, as well as the inter-modal ones. More recently, [182] employed the MAE framework to learn robust cross-modal representations. Their primary contribution is the MultiModal 3D Interaction module, which takes as input the concatenation of masked 3D volume features from both video frames and LiDAR sensors. Built upon self-attention layers and feed-forward networks, this module outputs a representation that is then split along the channel dimension and fed into modality-specific decoders to reconstruct the original inputs.
@@
3.11 Masking Strategy and Objective
A body of works have contributed to both the masking strategy and the objective function. These works either introduced a novel masking strategy that excels in hiding the salient information [1, 184, 40], or adapted the masking strategy to a different input type [122, 185]. Further, to maximize the benefit of using a new masking strategy, the objective function is altered to be more suitable for recovering the original input source [122, 39, 40, 186]. For example, [122] performed the masking in the discrete space of a dVAE and, due to this change, the reconstruction loss is the cross-entropy. Another example is the work of [185], in which the contrastive loss is employed to align the token representations of two different masked versions of the same image.
@@ 3.11 Masking Strategy and Objective
[186] slightly modified the MAE model by adding Gaussian noise to all pixels in the input image. With this change, besides reconstructing the masked regions, the objective is further extended to decode all denoised patches. The experiments demonstrated the superiority of this method over the original MAE. [40] introduced some improvements to the teacher-student MIM based on contrastive pre-training. The first addition is a ranking component that divides the patches into two subsets: the ones that contain salient information and the meaningless ones. The former subset is passed to the student, while the latter is given to the teacher, both being partially masked. The first objective is to reconstruct the masked patches of the student, as they are harder to predict due to containing more salient information. A second loss is used to align the globally encoded representations of the two subsets by maximizing the similarity between the embeddings of the [CLS] token. While the same encoder is used for both models, the gradients do not flow through the teacher.
@@
[5] introduced a masking strategy that selects important patches according to their similarity with a given text, computed in the CLIP embedding space. They further employed a contrastive loss to align text features with the features extracted from the unmasked patches.
@@
A couple of papers opted for a contrastive objective between two different views of an image obtained by distinct augmentations. Taking a different masking SSL approach, [187] proposed a method whose objective is to be more robust for downstream tasks. On the input image, two different sets of augmentations are applied. The augmented images are encoded, and a projection layer transforms the encoded representations into multivariate Gaussian distributions. Finally, a series of learnable masks are applied. The difference between both masked probabilistic embeddings is computed, the objective being that of minimizing this difference. [3] presented a pre-training framework that combines MIM and contrastive learning. Different subsets (a more aggressive one and a lighter one) of augmentations are applied to the input image to create two views, the latter having a portion of its tokens masked. Then, two vision transformer models (where the one for the unmasked image is an EMA of the other model) are used to encode the tokenized patches. The contrastive loss is applied only between the corresponding corrupted patches.
3.12 Masking Strategy and Task
@@
A few papers adopted MIM to increase the quality of images. The Saturated Mask AutoEncoder (SMAE), introduced by [193], is a two-stage approach for few-shot HDR imaging. The first stage focuses on representation learning, which is performed via masked image modeling. In this stage, the method creates two additional images from the original frame using exposure adjustment. Next, all three images are randomly masked with a high masking ratio before passing them to the model. LEMaRT [192] is an effective pre-training framework when applied on image harmonization as a downstream task. In this approach, the masked patches are replaced with the patches taken from a perturbed version of the original image. The authors also investigated what is the best strategy for creating the masks, concluding that random masking works best for image harmonization. Magnetic Resonance scans are generated in k-space and then transformed in the image domain with the inverse Fourier transform. To obtain an image with high quality and fidelity, the k-space needs to be fully sampled, but this is not realistic in most scenarios. To this end, [194] proposed to leverage the MAE framework by masking data in the k-space, represented in 3D: height, width and time, along the first dimension. By reconstructing the missing tokens via the $L^{1}$ loss, the adopted ViT models learn a rich feature representation that is able to estimate the unsampled data from k-space at inference. The estimated k-space is further refined using three sequential transformer-based decoders (one for each pair of dimensions), and employing the High-Dynamic Range loss after each decoder.
@@
3.13 Features and Task
@@
3.14 Architecture and Features
@@
3.15 Objective, Task and Theoretical Analysis
@@
3.16 Masking Strategy, Architecture and Objective
@@
3.17 Masking Strategy, Features and Objective
@@
3.18 Masking Strategy, Architecture, Task and Objective